# ReLaX-VQA: Residual Fragment and Layer Stack Extraction for Enhancing Video Quality Assessment

### v1

@@ 4.2 Performance comparison
    ReLaX-VQA: this version was separately trained and tested on each dataset;
@@
    ReLaX-VQA (w/o FT): this version was trained on LSVQ dataset and the frozen model was used for testing on the other datasets;
@@
    ReLaX-VQA (w/ FT): this version was trained on LSVQ dataset and the frozen model was fine-tuned on the other datasets to fit the new data features better.
@@
For ReLaX-VQA (w/o FT) and ReLaX-VQA (w/ FT) that were trained on LSVQ, we used the official subset LSVQtrain for training and the LSVQtest, another subset for testing to obtain frozen trained models. Furthermore, we selected the best-trained model based on $byrmse$ selection criteria and the results are reported in Table 4.2. We compare current state-of-the-art models using the correlation metrics SRCC and PLCC, including classical models and deep learning-based NR-VQA models. A first observation is that ReLaX-VQA (w/ FT) delivers the best performance across all datasets, followed by ReLaX-VQA. All ReLaX-VQA versions significantly outperform traditional NSS-based methods, such as BRISQUE. In comparison with the deep learning-based model VSFA, ReLaX-VQA (w/ FT) achieved $\Delta$ SRCC = 0.257 and $\Delta$ PLCC = 0.2982 on LIVE-VQC. ReLaX-VQA (w/o FT) achieves high scores of SRCC = 0.8312 and PLCC = 0.8427 on KoNViD-1k, not inferior to the best ones that are fine-tuned on the target dataset ( We need to note that for VSFA evaluation, we excluded the 110 2160p videos from the YouTube-UGC dataset because our GPU could not support them.). Most VQA methods achieve better performance on single-resolution datasets (i.e., KoNViD-1k), but this deteriorates on multi-resolution datasets (i.e., YouTube-UGC, LIVE-VQC). On the other hand, ReLaX-VQA maintains its high performance on datasets covering multiple resolutions. This indicates that ReLaX-VQA better captures the perception of quality in the vast video parameter space, indicating a wide adaptability and strong generalisation ability (see Table A.4 the Appendix for the full results).
4.3 Ablation studies
Table 3: Ablation study on spatio-temporal fragment sampling.
@@ 4.3 Ablation studies
Table 4: Ablation study on DNN layer stack.
@@
In our ablation study, we explored the influence of different components on the performance of our proposed method. In the first part (see Table 4), we compare the spatio-temporal features: frame difference and optical flow, along with their performance after residual fragment (RF) sampling and merged fragment (MF). We found that RF sampling of frame difference and optical flow can notably improve performance, especially when using ViT to extract frame difference features, with SRCC and PLCC improving by about 11.92% and 11.96%, respectively. The combination of ResNet-50 and ViT enhances performance by about 7.83% (SRCC) and 7.41% (PLCC) compared to using ResNet-50 alone, and by about 3.71% (SRCC) and 3.32% (PLCC) compared to using ViT alone to extract features. This result emphasizes the effectiveness of multiple DNN integration strategies in dealing with complex video content and the importance of preserving inter-frame spatio-temporal variations for video quality assessment.
In the second part (Table 4), we compare the layer stack and pool features of different DNNs. The results show that the layer-stacking framework captures complex video features more effectively compared to the pooling framework. In particular, the combination of ResNet-50 and ViT, layer stack + pool features, achieves SRCC of 0.7697 and PLCC of 0.7897, which is noticeably better than using only pooled features. This demonstrates the benefits of the layer-stacking framework in capturing and exploiting deep network features. Here, since ViT is based on the self-attention mechanism, we do not use a layer-stacking framework for this.
5 Conclusion and future work
In this paper, we proposed ReLaX-VQA to address the challenge of NR-VQA for UGC videos. We used the residual difference and optical flow fragments between successive sampled frames to capture the spatio-temporal information. Additionally, we enhanced DNN features by stacking layers to capture local and global information affecting video quality. We trained and tested our model on four public UGC datasets. Our results showed that ReLaX-VQA outperforms existing NR-VQA methods and demonstrates consistently high performance across the different datasets, achieving an overall average SRCC value of 0.8658 and PLCC value of 0.8872. Our method verified the benefit of residual information in enhancing the perceptual capability of NR-VQA models. Future research will focus on efficiency through the use of network architectures such as Swin Transformer or Mamba for feature extraction, as well as on the improvement of regression (e.g., KAN).
Acknowledgments and Disclosure of Funding
@@ 5 Conclusion and future work
Appendix A Appendix
A.1 Framework: DNN layer stack module
We stacked the extracted features using different layers of the deep neural network, fusing these multi-layered features to capture the quality-aware features. Finally, a simple MLP quality regression module maps the quality-aware features to video quality scores. Here, since ViT is based on the self-attention mechanism, we do not use a layer-stacking framework for this, shown in Fig. 2
Figure 2: The bespoke architectures of ResNet-50 Stack (I) and ResNet-50 Pool (II).
A.2 Spatio-temporal fragment sampling module
Here, we present the sampled frames and their corresponding ViT attention visualisations, along with the merged fragments (ours) and their ViT attention. These visualisations highlight the attention focus of the model. As shown in Fig. 3, the first video, $TelevisionClip\_1080P-68c6.mkv$, is a clip showing a man wearing an Angry Birds print moving in a snowy environment. The second video $Sports\_2160P-0455.mkv$ shows changes in perspective during the motions, including significant jittering. The videos can be downloaded and viewed at https://media.withyoutube.com/. Notice, that we get good results even when using our frozen ReLaX-VQA model without fine-tuning (ReLaX-VQA (w/o FT) to predict video quality scores.
Figure 3: Illustration of examples sampled from the YouTube-UGCwang2019youtube dataset. Top-row: $TelevisionClip\_1080P-68c6.mkv$. Bottom-row: $Sports\_2160P-0455.mkv$. GT refers to ground truth quality scores. Shown here are the scores predicted using ReLaX-VQA (w/o FT)
A.3 Nonlinear four-parameter logistic regression
In our experiments, PLCC and RMSE are calculated after a nonlinear four-parameter logistic regressionseshadrinathan2010study , where $\beta$ parameters were fitted using least squares.
$f(x)=\beta_{2}+\frac{\beta_{1}-\beta_{2}}{1+e^{-x+\frac{\beta_{3}}{|\beta_{4}|}}}\,,$ (15)
A.4 Performance comparison
We used the following datasets to train and evaluate our proposed model, CVD2014, KoNViD-1k, LIVE-VQC, YouTube-UGC and the largest dataset available, LSVQ. Based on the different strategies for training and testing we created three versions of our proposed method, ReLaX-VQA:
  1. i
    ReLaX-VQA: this version was separately trained and tested on each dataset;
  2. ii
    ReLaX-VQA (w/o FT): this version was trained on LSVQ dataset and the frozen model was used for testing on the other datasets;
  3. iii
    ReLaX-VQA (w/ FT): this version was trained on LSVQ dataset and the frozen model was fine-tuned on the other datasets to fit the new data features better.
We reported the SRCC, KRCC, PLCC, and RMSE metrics, along with the standard deviation (std) intervals, in Table A.4. We need to note that for VSFA evaluation, we excluded the 110 2160p videos from the YouTube-UGC dataset because our GPU could not support them. It is worth noting that for these metrics, we tested our frozen models using the same divisional test set as for the other NR-VQAs (ReLaX-VQA (w/o FT) and ReLaX-VQA (w/ FT)). Therefore, we performed only one test and the standard deviation intervals were not included. We included more results and tests in the supplementary materials.
Table 5: Full performance comparison of the evaluated NR-VQA models on the four NR-VQA datasets. The red and blue entries indicate the best and second-best performance on each database for each performance metric, respectively.
Datasets CVD2014nuutinen2016cvd2014
Models/Metrics SRCC(±STD) KRCC(±STD) PLCC(±STD) RMSE(±STD)
BRISQUEmittal2011blind 0.5553 (±0.0231) 0.3895 (±0.0161) 0.5527 (±0.0627) 18.4752 (±1.0274)
V-BLIINDS6705673 0.7306 (±0.0511) 0.5393 (±0.0473) 0.7853 (±0.0414) 13.7267 (±1.2130)
TLVQMkorhonen2019two 0.5399 (±0.0352) 0.4006 (±0.0298) 0.5785 (±0.0498) 18.0832 (±0.7052)
VIDEVALtu2021ugc 0.7663 (±0.0214) 0.5634 (±0.0221) 0.8062 (±0.0216) 13.1151 (±0.6579)
[…]
ReLaX-VQA (ours) 0.8643 (±0.0301) 0.6960 (±0.0442) 0.8895 (±0.0255) 9.8185 (±1.0969)
ReLaX-VQA (w/o FT) 0.7845 0.5930 0.8336 12.2445
ReLaX-VQA (w/ FT) 0.8974 0.7299 0.9294 8.1812
Datasets KoNViD-1khosu2017konstanz
Models/Metrics SRCC(±STD) KRCC(±STD) PLCC(±STD) RMSE(±STD)
BRISQUEmittal2011blind 0.6781 (±0.0083) 0.4935 (±0.0091) 0.6746 (±0.0069) 0.4753 (±0.0040)
V-BLIINDS6705673 0.7258 (±0.0115) 0.5322 (±0.0109) 0.7155 (±0.0107) 0.4498 (±0.0071)
TLVQMkorhonen2019two 0.7616 (±0.0067) 0.5635 (±0.0067) 0.7463 (±0.0065) 0.4285 (±0.0047)
VIDEVALtu2021ugc 0.8073 (±0.0174) 0.6036 (±0.0165) 0.7923 (±0.0169) 0.3928 (±0.0133)
[…]
ReLaX-VQA (ours) 0.8535 (±0.0218) 0.6594 (±0.0232) 0.8473 (±0.0214) 0.3370 (±0.0185)
ReLaX-VQA (w/o FT) 0.8312 0.6418 0.8427 0.3466
ReLaX-VQA (w/ FT) 0.8720 0.6881 0.8668 0.3211
Datasets LIVE-VQCsinno2018large
Models/Metrics SRCC(±STD) KRCC(±STD) PLCC(±STD) RMSE(±STD)
BRISQUEmittal2011blind 0.6096 (±0.0133) 0.4420 (±0.0127) 0.6652 (±0.0176) 12.7480 (±0.2697)
V-BLIINDS6705673 0.6925 (±0.0103) 0.5046 (±0.0111) 0.6872 (±0.0124) 12.4031 (±0.1948)
TLVQMkorhonen2019two 0.8133 (±0.0104) 0.6231 (±0.0105) 0.7912 (±0.0148) 10.4409 (±0.3123)
VIDEVALtu2021ugc 0.7725 (±0.0093) 0.5874 (±0.0080) 0.7752 (±0.0211) 10.7846 (±0.4082)
[…]
ReLaX-VQA (ours) 0.7655 (±0.0378) 0.5785 (±0.0365) 0.8079 (±0.0352) 9.8596 (±0.9099)
ReLaX-VQA (w/o FT) 0.7664 0.5812 0.8242 9.8201
ReLaX-VQA (w/ FT) 0.8468 0.6649 0.8876 7.9869
Datasets YouTube-UGCwang2019youtube
Models/Metrics SRCC(±STD) KRCC(±STD) PLCC(±STD) RMSE(±STD)
BRISQUEmittal2011blind 0.3517 (±0.0165) 0.2416 (±0.0119) 0.3768 (±0.0200) 0.6349 (±0.0056)
V-BLIINDS6705673 0.4783 (±0.0188) 0.3396 (±0.0143) 0.4967 (±0.0228) 0.5949 (±0.0094)
TLVQMkorhonen2019two 0.6802 (±0.0183) 0.4892 (±0.0168) 0.6876 (±0.0207) 0.4977 (±0.0132)
VIDEVALtu2021ugc 0.7814 (±0.0049) 0.5906 (±0.0052) 0.7929 (±0.0038) 0.4177 (±0.0033)
[…]
ReLaX-VQA (ours) 0.8014 (±0.0196) 0.6167 (±0.0193) 0.8204 (±0.0186) 0.3801 (±0.0177)
ReLaX-VQA (w/o FT) 0.8104 0.6131 0.8354 0.3768
ReLaX-VQA (w/ FT) 0.8469 0.6623 0.8652 0.3437

### v3

@@ IV-B Performance comparison
    ReLaX-VQA was trained and tested on each dataset (80% - 20% random split);
@@
    ReLaX-VQA (w/o FT) was trained on LSVQ, and the frozen model was used for testing on the other datasets;
@@
    ReLaX-VQA (w/ FT) was trained on LSVQ, and the frozen model was fine-tuned on the other datasets.
@@
For ReLaX-VQA (w/o FT) and ReLaX-VQA (w/ FT) that were trained on LSVQ, we used the official subset LSVQtrain for training and the LSVQtest to obtain frozen models. Furthermore, we selected the best-trained model based on $byrmse$ selection criteria and the results are reported in Table III-D. We compare current state-of-the-art (SOTA) models, including classical statistical and learning-based NR-VQA models, using correlation metrics SRCC, KRCC, PLCC, and RMSE. A key observation is that ReLaX-VQA (w/ FT) demonstrates superior performance across most datasets. Notably, ReLaX-VQA (w/ FT) achieves the highest correlation scores and RMSE values on three out of the four datasets (CVD2014, LIVE-VQC (for PLCC and RMSE), and YouTube-UGC). Furthermore, all versions of ReLaX-VQA significantly outperform traditional NSS-based methods, such as BRISQUE. Even without fine-tuning, ReLaX-VQA (w/o FT) delivers competitive results. For example, on the YouTube-UGC dataset, ReLaX-VQA (w/o FT) records a PLCC of 0.8354 and an RMSE of 0.3768, which are comparable to those of models that have been specifically fine-tuned on the target dataset. ReLaX-VQA (w/o FT) achieves high scores of SRCC = 0.8312 and PLCC = 0.8427 on KoNViD-1k, not inferior to the best ones that are fine-tuned on the target dataset ( We need to note that for the VSFA evaluation, we excluded the 110 2160p videos from the YouTube-UGC dataset because our GPU could not support them.). In comparison with the deep learning-based model VSFA, ReLaX-VQA (w/ FT) achieved $\Delta$ SRCC = 0.257 and $\Delta$ PLCC = 0.2982 on LIVE-VQC. Moreover, our model outperforms Fast-VQA-B, which also utilizes fragmentation, on two datasets, CVD2014 and YouTube-UGC. When fine-tuning is applied, ReLaX-VQA (w/ FT) achieves a $\Delta$ PLCC of 0.035 and a $\Delta$ RMSE of -0.0443 on the YouTube-UGC dataset. In comparison to the technical branch of DOVER, ReLaX-VQA (w/ FT) achieved $\Delta\text{SRCC}$ = 0.057 and $\Delta\text{PLCC}$ = 0.0303 on LIVE-VQC.
It is also worth noting that most of the SOTA methods achieve better performance on single-resolution datasets (i.e., KoNViD-1k), while ReLaX-VQA maintains its high performance on datasets covering multiple resolutions. This indicates that ReLaX-VQA better captures the perception of quality in the vast video parameter space, indicating a wide adaptability and strong generalisation ability.
IV-C Ablation studies, complexity analysis, and limitations
In our ablation study, we explored the influence of the different components on the performance of our proposed method. In the first part (see Table IV), we compare the spatio-temporal features: frame difference and optical flow, along with their performance after residual fragment (RF) sampling and merged fragment (MF). We found that RF sampling of frame difference and optical flow can notably improve performance, especially when using ViT to extract frame difference features, with SRCC and PLCC improving by about 11.92% and 11.96%, respectively. The combination of ResNet-50 and ViT enhances performance by about 7.83% (SRCC) and 7.41% (PLCC) compared to using ResNet-50 alone, and by about 3.71% (SRCC) and 3.32% (PLCC) compared to using ViT alone to extract features. This result emphasizes the effectiveness of multiple DNN integration strategies in dealing with complex video content and the importance of preserving inter-frame spatio-temporal variations for video quality assessment.
In the second part (Table V), we compare the layer stack and pool features of different DNNs. The results show that the layer-stacking framework captures complex video features more effectively compared to the pooling framework. In particular, the combination of ResNet-50 and ViT, ”layer stack + pool features”, achieves SRCC of 0.7697 and PLCC of 0.7897, which is noticeably better than using only pooled features. This demonstrates the benefits of the layer-stacking framework in capturing and exploiting deep network features. Here, since ViT is based on the self-attention mechanism, we do not use a layer-stacking framework for this.
TABLE III: FLOPs and running time (on CPU, average of ten runs) comparison of ReLaX-VQA and other deep learning SOTA methods across different spatial resolutions.
Method 540P 720P 1080P
FLOPs (G) Time (s) FLOPs (G) Time (s) FLOPs (G) Time (s)
VSFA [7] 10249 255.59 18184 466.88 40919 1004.02
FAST-VQA-B [14] 279 24.70 279 23.81 279 24.74
DOVER [44] 282 11.95 292 12.02 282 12.48
ReLaX-VQA (ours) 927 31.08 927 35.93 927 47.08
To ensure a fair comparison of the computational efficiency of different methods, all tests were conducted on the same workstation equipped with an NVIDIA GeForce RTX 3090 GPU, a 6-core Intel(R) Xeon(R) W-1250 CPU, and 64 GB of RAM. We used the same video from the KoNViD-1k database for testing at various resolutions. Each test was repeated 10 times, and the average computation time (in seconds) for each method was reported. Table III compares the three best-performing deep learning-based methods in terms of their computational and runtime complexity (CPU)  [^1] Since part of our algorithm (frame differencing, optical flow estimation, and fragment ranking) cannot be fully accelerated on the GPU, including a GPU time comparison would not provide a fair assessment.. It is evident that ReLaX-VQA, Fast-VQA-B and DOVER maintain stable FLOP numbers across resolutions, resulting in small differences in runtime. For example, ReLaX-VQA requires 31.08 seconds at 540P and 47.08 seconds at 1080P. In contrast, VSFA exhibits a significant increase in both FLOPs and runtime as resolution increases, showing a notably longer runtime on the CPU. While FAST-VQA-B and DOVER are faster, ReLaX-VQA achieves a good balance, maintaining a reasonable computational cost while demonstrating consistent efficiency and better generalization.
One limitation of ReLaX-VQA is the video frame sampling strategy. We sampled two frames every half-second to obtain successive, temporally correlated frame pairs. However, the extent of correlation depends on camera/scene motion, meaning sampled frames have a small chance of not being able to capture the content changes if the camera motion is stationary. Future work will attempt to further optimize this strategy by taking into account the hierarchical encoding of frames within the Group Of Pictures (GOP).
TABLE IV: Ablation study on spatio-temporal fragment sampling.
@@ IV-C Ablation studies, complexity analysis, and limitations
TABLE V: Ablation study on DNN layer stack.
@@
V Conclusion and future work
In this paper, we propose ReLaX-VQA to address the challenges of NR-VQA for UGC videos. We employed residual difference and optical flow fragments between successive sampled frames along with a ranking process to capture the spatio-temporal information. Additionally, we enhanced DNN features by stacking layers to capture local and global information affecting video quality. We trained and tested our model on four public UGC datasets. Our results showed that ReLaX-VQA outperforms existing NR-VQA methods and demonstrates consistently high performance across the different datasets, achieving an overall average SRCC value of 0.8658 and PLCC value of 0.8873. Our method verified the benefit of residual information in enhancing the perceptual capability of NR-VQA models. Future research will focus on improving efficiency through network architectures (e.g., Swin Transformer or Mamba for feature extraction) and enhancing regression approaches (e.g., KAN) to optimise memory usage and processing speed.
Acknowledgment